1.      Find the sample size to estimate a population mean within 95% confidence if the standard deviation is 6.2.


1
Expert's answer
2022-04-07T05:50:58-0400

"Z_{95}=1.96"

"n=(\frac{Z_{95}\sigma}{0.5})^2=(\frac{1.96x6.2}{0.5})^2=590.7"

0.5 is a valid error. Usually 50%=0.5 is used, unless otherwise specified in the task.